OpenStudy (anonymous):

How do i do this? Solve –3(2x – 1) = x + 10. You must show all work.

OpenStudy (anonymous):

-6x+3=x+10 -7x=7 x=-1

OpenStudy (angela210793):

-6x+3-x-10=0 -7x=7 x=-1